# SMOGAN: Synthetic Minority Oversampling with GAN Refinement for Imbalanced Regression

### v1

@@ Abstract
Imbalanced regression refers to prediction tasks where the target variable is skewed. This skewness hinders machine learning models, especially neural networks, which concentrate on dense regions and therefore perform poorly on underrepresented (minority) samples. Despite the importance of this problem, only a few methods have been proposed for imbalanced regression. Many of the available solutions for imbalanced regression adapt techniques from the class imbalance domain, such as linear interpolation and the addition of Gaussian noise, to create synthetic data in sparse regions. However, in many cases, the underlying distribution of the data is complex and non-linear. Consequently, these approaches generate synthetic samples that do not accurately represent the true feature–target relationship. To overcome these limitations, we propose SMOGAN, a two-step oversampling framework for imbalanced regression. In Stage 1, an existing oversampler generates initial synthetic samples in sparse target regions. In Stage 2, we introduce DistGAN, a distribution-aware GAN that serves as SMOGAN’s filtering layer and refines these samples via adversarial loss augmented with a Maximum Mean Discrepancy objective, aligning them with the true joint feature–target distribution. Extensive experiments on 23 imbalanced datasets show that SMOGAN consistently outperforms the default oversampling method without the DistGAN filtering layer.
[…]
123
@@ 1 Introduction
Imbalanced data in regression tasks that predict continuous outcomes can compromise model performance [36]. In such settings, some target ranges, particularly minority samples, have very few observations [8, 55]. Standard regression models tend to focus on densely populated regions of the target space, resulting in poor predictions for underrepresented values [17, 8]. Yet, accurately forecasting these uncommon targets is essential for many practical applications [55].
Conversely, in both binary and multi-class classification, imbalanced data occurs when some classes have far fewer samples than others. This leads classifiers to focus on the majority classes while missing the sparse patterns of minority classes. Effective techniques for addressing class imbalance are essential to ensure that all classes, even those with fewer examples, are accurately recognized [30, 16, 12, 40]. Critical applications like fraud detection, medical diagnostics, and fault detection are particularly affected, as important minority classes might be missed [28, 35].
In various critical domains, imbalanced regression directly addresses the challenge of predicting rare, high-impact events amid a sea of routine observations. For instance, in economics, models must accurately forecast rare financial downturns against generally stable market trends. In meteorology, imbalanced regression is vital for predicting extreme weather events that can have devastating impacts despite their infrequency. Similarly, in crisis management, the ability to anticipate unlikely emergencies can be crucial for timely responses, while in fault diagnosis, detecting rare anomalies in complex systems is essential for ensuring operational safety and reliability. These examples are just a few illustrations of the broader applicability of imbalanced regression techniques, which extend across many other fields and real-world scenarios [36].
@@
Existing oversampling methods for imbalanced regression often oversimplify the problem and overlook the complex underlying structure of continuous target distributions. To address this limitation, we propose SMOGAN, a two-step framework that integrates DistGAN (a distribution-aware generative adversarial network [26]) as an inbuilt filtering layer to refine synthetic samples and preserve real data fidelity. As illustrated in Figure 2, in Stage 1 (Initial Synthetic Generation), a base oversampler (e.g, SMOGN) produces initial synthetic samples in sparse target regions. In Stage 2 (Adversarial Refinement), these candidates are passed through DistGAN’s generator and discriminator: the generator refines them by optimizing a combined adversarial loss augmented with Maximum Mean Discrepancy (MMD) to minimize kernel-based distribution discrepancies, while the discriminator, trained exclusively on real minority samples, filters out-of-distribution outputs. We conduct comprehensive experiments on 23 benchmark imbalanced regression datasets. We extend our evaluation across diverse application domains to illustrate the effectiveness and generalizability of the proposed SMOGAN framework. Importantly, SMOGAN’s modular design allows the initial oversampler in Stage 1 to be replaced with any data-level technique of the user’s choice, enabling seamless integration of new or domain-specific methods with DistGAN’s refinement process.
@@ 2.1 Class imbalance
Chawla et al. [16] introduced the first data-level class imbalance handling approach, SMOTE, which generated synthetic minority samples $\tilde{\mathbf{x}}$ by interpolating between a minority instance $\mathbf{x}_{i}$ and one of its $k$ nearest neighbours $\mathbf{x}_{i,\mathrm{nn}}$ according to
@@
SMOTE became the de facto method for handling class imbalance, with many extensions proposed thereafter. Han et al. [29] proposed Borderline-SMOTE, which oversampled only those minority instances near the class boundary. Bunkhumpornpat et al. [13] presented DBSMOTE, guiding interpolation toward high-density minority regions. Many other SMOTE variations were proposed for the class imbalance problem [7, 6, 31, 19, 57, 20, 21, 25, 50, 18, 41, 48, 38, 43].
Along with SMOTE and its variations, advanced techniques such as generative adversarial networks (GANs) were extensively explored in the class imbalance literature. Mariani et al. [42] developed BAGAN, a class-conditional GAN initialized with an autoencoder to produce diverse, high-quality minority images. Tanaka and Aranha [52] demonstrated that GAN-generated tabular data could replace real samples in classifier training and improve minority recall. Engelmann and Lessmann [24] designed a conditional WGAN with gradient penalty, auxiliary classifier loss, Gumbel-softmax for categorical features, and cross-layer interactions for mixed-type tables. Sharma et al. [49] embedded SMOTE-interpolated points into a GAN generator to diversify synthetic minority examples for addressing class imbalance. Jiang et al. [34] employed an encoder–decoder–encoder GAN on time-series data to detect anomalies via combined reconstruction and adversarial losses. Lee and Park [39] generated rare-attack intrusion samples with a GAN and retrained a random forest to boost detection.
@@ 2.2 Imbalanced regression
Unlike class imbalance, the imbalanced regression problem received relatively little attention. Torgo et al. [55] and Torgo and Ribeiro [53] introduced the first data-level method (SMOTER) by extending SMOTE to regression with a relevance function $\phi(y)$ to identify rare targets, undersample common values, and interpolate rare points. Branco et al. [9] generalized SMOTER by adding Gaussian noise (SMOGN) to increase sample variability, and later proposed WERCS [10], which probabilistically oversampled or undersampled based on $\phi(y)$ and simple random or noise generation. Alahyari and Domaratzki [1] proposed LDAO, a local distribution–based adaptive oversampling method that clustered the joint feature–target space and drew synthetic samples per cluster to balance the data. Camacho et al. [15] introduced G-SMOTE, incorporating target-space geometry into interpolation, and Camacho and Bacao [14] replaced hard rarity thresholds with instance-based weighting in WSMOTER. Moniz et al. [44] integrated SMOTE oversampling into boosting (SMOTEBoost), demonstrating the adaptability of data-level strategies for regression imbalance.
Alongside data-level strategies, several algorithmic methods were proposed for imbalanced regression. Cost-sensitive learning extended standard regression by weighting errors on rare targets [61, 23, 22]. Steininger et al. [51] introduced DenseLoss, which emphasized low-density targets in the loss function. Yang et al. [60] applied label-distribution smoothing in deep models. Ren et al. [45] proposed Balanced MSE to scale error contributions according to target rarity. Ribeiro and Moniz [46] developed SERA, a squared-error relevance-weighted metric for evaluation.
@@
GAN-based methods remained largely unexplored in the imbalanced regression domain. Hssayeni et al. [32] proposed a convolutional conditional GAN that predicted continuous targets from time-series inputs, demonstrating improved generalization. Jia et al. [33] introduced RNGAN, which employed adversarial gating of regression and neighbour-based features and assessed sample fidelity via a correlation-change metric. More robust advanced techniques similar to those in class imbalance research merited exploration to generate high-quality synthetic data that respected the underlying distribution of real data, particularly in sparse regions.
3 SMOGAN method
@@ 3.1 Stage 1: Initial Synthetic Generation
We base our initial oversampling on SMOGN (SmoteR and Gaussian noise) [9], the current state-of-the-art regression oversampler, chosen for its ability to both interpolate and inject variability in sparse target regions. We initialize by generating an initial pool of synthetic samples using SMOGN. Each target $y$ in the original data set is assigned a relevance score $\phi(y)\in[0,1]$, labeling values within the interquartile range as unimportant and those in the tails as increasingly important via a simple sigmoid function. We then split the data into rare ($\phi(y)\geq t_{R}$) and normal ($\phi(y)<t_{R}$) cases, using a relevance threshold $t_{R}$ (typically 0.8).
Focusing only on the rare set, each seed sample $(\mathbf{x}_{i},y_{i})$ spawns synthetic neighbours as follows: 1. Select one of its $k$ nearest rare neighbours $(\mathbf{x}_{j},y_{j})$ and compute the Euclidean distance $d_{ij}=\|\mathbf{x}_{j}-\mathbf{x}_{i}\|$. 2. Let $d^{\star}$ be half the median of these $k$ distances. If $d_{ij}<d^{\star}$, interpolate (SmoteR):
@@ 3.2 Stage 2: Adversarial Refinement with DistGAN
Having generated the initial synthetic pool $\mathcal{D}_{\mathrm{syn}}^{(0)}$, we refine it by training a Wasserstein GAN with gradient penalty (WGAN-GP) [4, 27]. Unlike a standard GAN, where the discriminator is trained as a binary classifier, the WGAN framework replaces it with a critic that scores sample realness directly and optimizes the Wasserstein (Earth-Mover) distance between distributions. In the GP variant, weight clipping is replaced by a gradient-penalty term that enforces the $1$-Lipschitz constraint, yielding smoother gradients and more stable training. On top of this, the generator loss incorporates an RBF-kernel MMD term to align all moments between refined and real minority samples. Figure 3 illustrates this two-stage adversarial refinement workflow.
@@ 3.2.1 Critic objective
$\mathcal{L}_{D}$ $=\underbrace{\mathbb{E}_{x\sim\mathcal{D}_{\mathrm{syn}}^{(0)}}\!\bigl{[}D(G(x))\bigr{]}-\mathbb{E}_{z\sim\mathcal{D}_{\mathrm{real}}}\!\bigl{[}D(z)\bigr{]}}_{\text{Wasserstein objective}}$
$\quad+\,\lambda_{\mathrm{gp}}\;\underbrace{\mathbb{E}_{\begin{subarray}{c}z,x,\epsilon\end{subarray}}\bigl{(}\|\nabla_{\hat{u}}D(\hat{u})\|_{2}-1\bigr{)}^{2}}_{\text{gradient penalty}}.$
We set $\lambda_{\mathrm{gp}}=10$, following the hyper-parameter recommendation in the original WGAN-GP paper  [27].
@@ 3.2.2 Generator objective
$\mathcal{L}_{G}$ $=\underbrace{-\mathbb{E}_{x\sim\mathcal{D}_{\mathrm{syn}}^{(0)}}D\bigl{(}G(x)\bigr{)}}_{\text{adversarial term}}$
$\quad+\,\alpha\;\underbrace{\widehat{\mathrm{MMD}}^{2}\bigl{(}\{G(x_{i})\},\{z_{j}\}\bigr{)}}_{\text{moment\textendash matching term}},$
@@ 4 Evaluation methodology
We benchmark SMOGAN against SMOGN, the current state-of-the-art imbalanced regression oversampler that extends SMOTER with Gaussian noise, using a diverse collection of benchmark datasets. SMOGN requires specifying a relevance function $\phi(y)$ and a relevance threshold $t_{R}$; in our experiments, we employ the default threshold $t_{R}=0.8$, as recommended by the authors across all datasets. First, we evaluate SMOGN with its default parameter settings. Then we integrate the DistGAN layer into SMOGN to form SMOGAN, with the primary goal of refining the synthetic sample pool and demonstrating improved sample quality under this configuration. For reference, we also include a baseline model trained without any oversampling, but because SMOGN (and thus SMOGAN) can and in practice should be fine-tuned to specific domains, our main comparison focuses on the default versions of SMOGN versus SMOGAN, with baseline results provided solely for illustrative purposes.
@@ 4.1 Datasets
We evaluated our method using 23 datasets from three sources: the Keel repository [2], the DataSets-IR collection [10], and the imbalancedRegression repository [5]. These datasets span multiple domains and vary in size, dimensionality, and degree of imbalance, making them benchmarks for the rigorous evaluation of imbalanced regression methods and enabling fair comparisons. Table 1 shows the number of instances and features for each dataset.

### v2

@@ Abstract
Imbalanced regression refers to prediction tasks where the target variable is skewed. This skewness hinders machine learning models, especially neural networks, which concentrate on dense regions and therefore perform poorly on underrepresented (minority) samples. Despite the importance of this problem, only a few methods have been proposed for imbalanced regression. Many of the available solutions for imbalanced regression adapt techniques from the class imbalance domain, such as linear interpolation and the addition of Gaussian noise, to create synthetic data in sparse regions. However, in many cases, the underlying distribution of the data is complex and non-linear. Consequently, these approaches generate synthetic samples that do not accurately represent the true feature-target relationship. To overcome these limitations, we propose SMOGAN, a two-step oversampling framework for imbalanced regression. In Stage 1, an existing oversampler generates initial synthetic samples in sparse target regions. In Stage 2, we introduce DistGAN, a distribution-aware GAN that serves as SMOGAN’s filtering layer and refines these samples via adversarial loss augmented with a Maximum Mean Discrepancy objective, aligning them with the true joint feature-target distribution. Extensive experiments on 23 imbalanced datasets show that SMOGAN consistently outperforms the default oversampling method without the DistGAN filtering layer.
Keywords: GAN, Imbalanced, Regression, Generative AI, Unsupervised learning
@@ 1 Introduction
Imbalanced data in regression tasks that predict continuous outcomes can compromise model performance (Krawczyk, 2016). In such settings, some target ranges, particularly minority samples, have very few observations (Branco et al., 2016; Torgo et al., 2013). Standard regression models tend to focus on densely populated regions of the target space, resulting in poor predictions for underrepresented values (Chawla et al., 2004; Branco et al., 2016). Yet, accurately forecasting these uncommon targets is essential for many practical applications (Torgo et al., 2013).
Conversely, in both binary and multi-class classification, imbalanced data occurs when some classes have far fewer samples than others. This leads classifiers to focus on the majority classes while missing the sparse patterns of minority classes. Effective techniques for addressing class imbalance are essential to ensure that all classes, even those with fewer examples, are accurately recognized (He and Garcia, 2009; Chawla et al., 2002; Buda et al., 2018; Liu et al., 2009). Critical applications like fraud detection, medical diagnostics, and fault detection are particularly affected, as important minority classes might be missed (Guo et al., 2017; Johnson and Khoshgoftaar, 2019).
Imbalanced regression impacts many critical real-world applications where accurately predicting rare but high-impact events in the tails of skewed distributions is paramount. For instance, In air quality forecasting, infrequent high ozone concentration events are systematically underestimated by SVR models trained on abundant moderate values, posing health risk underpredictions (Zhen et al., 2025). In intensive care, predicting prolonged ventilation days from skewed ICU datasets is hampered by scarce samples of long-duration cases, leading to suboptimal resource planning (Radovanović et al., 2025). In precision agriculture, vineyard yield datasets feature extreme low- and high-yield blocks that are poorly represented, prompting conditional UNet–ConvLSTM architectures with zonal weighting to address these extremes (Kamangir et al., 2024). In chemical synthesis planning, high-yield reactions are rarer yet more valuable, yet existing prediction models overlook this imbalance and underperform on critical high-yield bins (Ma et al., 2024). These cases underscore the inadequacy of conventional regression and the need for specialized imbalanced-regression methodologies.
@@
Existing oversampling methods for imbalanced regression often oversimplify the problem and overlook the complex underlying structure of continuous target distributions. To address this limitation, we propose SMOGAN, a two-step framework that integrates DistGAN (a distribution-aware generative adversarial network (Goodfellow et al., 2014)) as an inbuilt filtering layer to refine synthetic samples and preserve real data fidelity. As illustrated in Figure 2, in Stage 1 (Initial Synthetic Generation), a base oversampler (e.g., SMOGN) produces initial synthetic samples in sparse target regions. In Stage 2 (Adversarial Refinement), these candidates are passed through DistGAN’s generator and discriminator: the generator refines them by optimizing a combined adversarial loss augmented with Maximum Mean Discrepancy (MMD) to minimize kernel-based distribution discrepancies, while the discriminator, trained exclusively on real minority samples, filters out-of-distribution outputs. We conduct comprehensive experiments on 23 benchmark imbalanced regression datasets. We extend our evaluation across diverse application domains to illustrate the effectiveness and generalizability of the proposed SMOGAN framework. Importantly, SMOGAN’s modular design allows the initial oversampler in Stage 1 to be replaced with any data-level technique of the user’s choice, enabling seamless integration of new or domain-specific methods with DistGAN’s refinement process.
@@ 2.1 Class imbalance
This section reviews key methods for addressing class imbalance in classification, as these techniques serve as the basis for several regression approaches we evaluate. The foundational work by Chawla et al. (2002) presented SMOTE (Synthetic Minority Over-sampling Technique), the first data-level solution to class imbalance. SMOTE creates artificial minority class examples $\tilde{\mathbf{x}}$ through linear interpolation between a minority sample $\mathbf{x}_{i}$ and one of its $k$ nearest minority neighbors $\mathbf{x}_{i,\mathrm{nn}}$:
@@
This method established itself as the standard approach for managing imbalanced datasets, inspiring numerous extensions and modifications. The core interpolation mechanism of SMOTE provides crucial insights for understanding its regression counterparts, particularly SMOTER and SMOGN, which we analyze in our experiments (Torgo et al., 2013; Branco et al., 2017).
Beyond SMOTE and its derivatives, researchers have applied generative adversarial networks (GANs) to tackle class imbalance problems. Mariani et al. (2018) introduced BAGAN, which combines an autoencoder-based initialization with class-conditional generation to create varied and realistic minority class images. In the tabular domain, Tanaka and Aranha (2019) showed that synthetic data from GANs could effectively substitute real training samples while enhancing minority class detection rates. For mixed-type tabular data, Engelmann and Lessmann (2020) proposed a conditional WGAN architecture incorporating gradient penalties, auxiliary classification objectives, Gumbel-softmax activation for discrete variables, and cross-layer connections.
Recent work has also integrated traditional oversampling with deep generative models. Sharma et al. (2022) combined SMOTE’s interpolation strategy with GAN-based generation, using interpolated samples to guide the generator toward more diverse synthetic minority instances to tackle class imbalance. Applications extend to sequential data, where Jiang et al. (2019) developed an encoder-decoder-encoder GAN framework for time-series anomaly detection using joint reconstruction and adversarial objectives. Similarly, Lee and Park (2021) addressed rare attack detection in network intrusion systems by generating synthetic attack patterns with GANs and subsequently retraining random forest classifiers.
@@ 2.2 Imbalanced regression
Imbalanced regression has received far less attention than classification imbalance. Torgo and Ribeiro (2007) and Torgo et al. (2013) developed SMOTER, the first data-level method for regression. SMOTER uses a relevance function $\phi(y)$ to find rare target values, then undersamples common cases and interpolates between rare ones.
Branco et al. (2017) improved SMOTER by adding Gaussian noise (SMOGN) to create more varied synthetic samples. They later developed WERCS (Branco et al., 2019), which uses $\phi(y)$ to decide whether to oversample or undersample each instance using random selection or noise. Camacho et al. (2022) developed G-SMOTE, which considers the geometry of target values when interpolating. Camacho and Bacao (2024) replaced SMOTER’s fixed thresholds with continuous weights in WSMOTER. Alahyari and Domaratzki (2025) proposed LDAO, which clusters the joint feature-target space and applies kernel density estimation within each cluster to generate synthetic samples that preserve local distribution characteristics. Alahyari et al. (2025) used Mahalanobis-GMM detection with GAN-based generation for imbalanced regression, eliminating manual threshold selection while preserving statistical properties of minority samples. Moniz et al. (2018) combined SMOTE with boosting in SMOTEBoost, showing that oversampling works well with ensemble methods for regression.
Researchers have also developed algorithmic methods for imbalanced regression. Cost-sensitive learning weights errors on rare targets more heavily than common ones (Zhou and Liu, 2010; Elkan, 2001; Domingos, 1999). Steininger et al. (2021) created DenseLoss, which gives more importance to low-density targets in the loss function. Yang et al. (2021) used label-distribution smoothing in neural networks to handle imbalance. Ren et al. (2022) introduced Balanced MSE, which scales errors based on how rare each target value is.
New evaluation metrics have been developed specifically for imbalanced regression. Ribeiro and Moniz (2020) created SERA (Squared Error Relevance Area), which weights errors based on target rarity. Unlike RMSE, which treats all errors equally, SERA gives more weight to errors on rare values by integrating across different relevance thresholds. Standard metrics hide poor performance on rare samples because errors on common samples dominate the overall score. This makes metrics like SERA essential for evaluating imbalanced regression methods properly.
3 SMOGAN method
@@ 3.1 Stage 1: Initial Synthetic Generation
We base our initial oversampling on SMOGN (SmoteR and Gaussian noise) (Branco et al., 2017), a popular regression oversampler, chosen for its ability to both interpolate and inject variability in sparse target regions. We initialize by generating an initial pool of synthetic samples using SMOGN. Each target $y$ in the original data set is assigned a relevance score $\phi(y)\in[0,1]$, labeling values within the interquartile range as unimportant and those in the tails as increasingly important via a simple sigmoid function. We then split the data into rare ($\phi(y)\geq t_{R}$) and normal ($\phi(y)<t_{R}$) cases, using a relevance threshold $t_{R}$ (typically 0.8).
Focusing only on the rare set, each seed sample $(\mathbf{x}_{i},y_{i})$ spawns synthetic neighbours as follows: 1. Select one of its $k$ nearest rare neighbours $(\mathbf{x}_{j},y_{j})$ and compute the Euclidean distance $d_{ij}=\|\mathbf{x}_{j}-\mathbf{x}_{i}\|$. 2. Let $d^{\star}$ be half the median of $k$ distance. If $d_{ij}<d^{\star}$, interpolate (SmoteR):
@@ 3.2 Stage 2: Adversarial Refinement with DistGAN
Having generated the initial synthetic pool $\mathcal{D}_{\mathrm{syn}}^{(0)}$, we refine it by training a Wasserstein GAN with gradient penalty (WGAN-GP) (Arjovsky et al., 2017; Gulrajani et al., 2017). We choose WGAN-GP because it provides stable training dynamics and generates diverse, high-quality samples without mode collapse, critical for preserving the complex structure of minority regions in tabular data. In the GP variant, weight clipping is replaced by a gradient-penalty term that enforces the $1$-Lipschitz constraint, yielding smoother gradients and more stable training. On top of this, the generator loss incorporates an RBF-kernel MMD term to align all moments between refined and real minority samples. Figure 3 illustrates this two-stage adversarial refinement workflow.
@@ 3.2.1 Critic objective
$\mathcal{L}_{D}$ $=\underbrace{\mathbb{E}_{x\sim\mathcal{D}_{\mathrm{syn}}^{(0)}}\!\bigl{[}D(G(x))\bigr{]}-\mathbb{E}_{z\sim\mathcal{D}_{\mathrm{real}}}\!\bigl{[}D(z)\bigr{]}}_{\text{Wasserstein objective}}+\,\lambda_{\mathrm{gp}}\;\underbrace{\mathbb{E}_{\begin{subarray}{c}z,x,\epsilon\end{subarray}}\bigl{(}\|\nabla_{\hat{u}}D(\hat{u})\|_{2}-1\bigr{)}^{2}}_{\text{gradient penalty}}.$
We set $\lambda_{\mathrm{gp}}=10$, following the hyper-parameter recommendation in the original WGAN-GP paper  (Gulrajani et al., 2017).
@@ 3.2.2 Generator objective
$\mathcal{L}_{G}$ $=\underbrace{-\mathbb{E}_{x\sim\mathcal{D}_{\mathrm{syn}}^{(0)}}D\bigl{(}G(x)\bigr{)}}_{\text{adversarial term}}+\,\alpha\;\underbrace{\widehat{\mathrm{MMD}}^{2}\bigl{(}\{G(x_{i})\},\{z_{j}\}\bigr{)}}_{\text{moment-matching term}},$
@@ 4 Evaluation methodology
We benchmark SMOGAN against SMOGN, the current state-of-the-art imbalanced regression oversampler that extends SMOTER with Gaussian noise, using a diverse collection of benchmark datasets. SMOGN requires specifying a relevance function $\phi(y)$ and a relevance threshold $t_{R}$; in our experiments, we employ the default threshold $t_{R}=0.8$, as recommended by the authors across all datasets. First, we evaluate SMOGN with its default parameter settings. Then we integrate the DistGAN layer into SMOGN to form SMOGAN, with the primary goal of refining the synthetic sample pool and demonstrating improved sample quality under this configuration. We further evaluate SMOGAN against a GAN-only approach that skips the initial SMOGN oversampling step. We also compare it to a baseline with no oversampling and to other state-of-the-art methods, including random oversampling (RO), G-SMOTE, and WSMOTER.
@@ 4.1 Datasets
We evaluated our method using 23 datasets from three sources: the Keel repository (Alcalá-Fdez et al., 2011), the DataSets-IR collection (Branco et al., 2019), and the imbalancedRegression repository (Avelino et al., 2024). These datasets span multiple domains and vary in size, dimensionality, and degree of imbalance, making them benchmarks for the rigorous evaluation of imbalanced regression methods and enabling fair comparisons. Table 1 shows the number of instances and features for each dataset.